# Hyper-Compression: Model Compression via Hyperfunction

### v1

@@ VI. Experimental Details
  • Details of experiments related to UNet: We perform parameter compression on both UNet and the pruned UNet, and test the performance loss of the models after compression on the dataset Carvana [1].
@@
      Pruned UNet: Specifically, the pruned UNet refers to the UNet model that has undergone Taylor-Rank Pruning [16] with FLOPs Regularization [11] (strength=0.001). This is essentially a channel pruning method. The bar chart below illustrates the percentage of channels pruned in each layer of the pruned UNet relative to the original UNet.
@@
      Evaluation: The Dice coefficient, also known as the Dice similarity index, is a statistical measure used to gauge the similarity between two sets. It is particularly popular in the field of image segmentation to assess the accuracy of models by comparing the overlap between the predicted segmentation and the ground truth. Mathematically, the Dice coefficient is defined as
@@
      where $A$ represents the predicted segmentation, and $B$ is the ground truth. The value of the Dice coefficient ranges from 00 to $1$, where $1$ indicates perfect agreement between two sets, and 00 indicates no overlap at all.
@@
      MobileNetV3: MobileNetV3 is a state-of-the-art convolutional neural network architecture designed specifically for mobile devices. Building upon the success of its predecessors, MobileNetV3 introduces novel design principles and optimization techniques to achieve a compelling balance between accuracy and efficiency.
@@
      Dataset: The CIFAR-10 dataset [14] is a well-known benchmark in the field of machine learning and computer vision, consisting of $6,000$ $32\times 32$ colored images belonging to $10$ classes. The dataset is split into $50,000$ training images and $10,000$ testing images.
@@
  • Details of experiments related to LlaMA2-7B: We apply our compression method to the LLaMA2-7B [21] and three other performant pruned variants of the LLaMA2-7B: Sheared-LLaMA-1.3B [25], TinyLLaMA [27], and LiteLLaMA. We then conduct performance tests on these models using various metrics, including SciQ [23], WinoGrande [19], ARC-E [5], 25-shot ARC-C [5], 10-shot Hellaswag [26], 32-shot BoolQ [4], 32-shot NQ [15], 5-shot MMLU [12], and Perplexity [24] to evaluate the effects of compression by using the package lm-evaluation-harness [9].
@@
      Evaluation: In this study, most evaluation metrics for large language models (LLMs) are tested using the open-source package lm-evaluation-harness [9] which is a versatile and open-source Python package for benchmarking large language models (LLMs) across various tasks and datasets. The "(i)" after the name of an evaluation metric means "i-shot".
@@ 5.1 Compression Performance
Model PPL
LlaMA2-7B 5.19
LlaMA2-7B + HF 5.48 (+0.29)
Sheared-LlaMA-1.3B 7.76
Sheared-LlaMA-1.3B + HF 7.99 (+0.23)
TinyLlaMA 7.32
TinyLlaMA + HF 7.54 (+0.22)
LiteLlaMA 20.22
LiteLlaMA + HF 23.89 (+3.67)
Figure 11: An example output from LlaMA2-7B, Sheared-LlaMA-1.3B, TinyLlaMA, LiteLlaMA and their compressed model by using our method. These results demonstrate that our compression method effectively preserves the models’ ability to produce meaningful and grammatically correct text.
Here, we show the compression performance of LLaMA2-7B and seven different variants based on LlaMA2-7B. First, in Table 4, we test eight important downstream tasks. By directly compressing LLaMA2-7B using our method (LLaMA2-7B + HF), a compression ratio of 2.6$\times$ is achieved with only a 0.99% decrease in the average score. When our method is combined with other LLaMA2-7B variants (Sheared-LlaMA-1.3B, TinyLlaMA, and LiteLlaMA) respectively, higher compression ratios can be achieved, with the change in the average score remaining within 1%, compared to the pruned ones. Moreover, Perplexity (PPL) is an important metric for evaluating the overall performance of large language models (LLMs). We conduct tests on the publicly available dataset wikitext-2-raw-v1  [^2] For more details, please visit the website of dataset wikitext-2-raw-v1: https://huggingface.co/datasets/Salesforce/wikitext.. As shown in Table 5, the PPL of LLaMA2-7B+ET increases by only 0.29 compared to the original LLaMA2-7B model, whereas the PPL increases caused by the other three LLaMA2-7B variants (Sheared-LlaMA-1.3B, TinyLlaMA, LiteLlaMA) are 2.57, 2.13, and 15.03, respectively. This demonstrates that our method not only enhances the compression ratio but also effectively maintains the model’s overall performance. Figure 11 presents two examples comparing the text outputs generated by eight different LLM models for a given prompt. The experimental data demonstrate that our method achieves an preferable compression ratio while having minimal impact on the model’s performance across various downstream tasks. Moreover, it can also integrate effectively with other compression methods.
@@ 5.2 Compression Time
As shown in Table 6, we record the compression times for eight models using our method. The time to compress smaller models like UNet and MobileNetV3 is only 35 seconds, with Pruned-MobileNetV3 requiring only 8.83 seconds. The compression time for large language models is longer. For example, compressing LLaMA2-7B takes 53 minutes, while LiteLLaMA is compressed in just 6 minutes. The smaller the model’s size, the faster the compression is.
@@ 5.3 Inference Time
We test the inference time and the execution time of our method under different batch sizes on NVIDIA GeForce RTX 4060 and NVIDIA A6000 graphics cards, respectively.
@@
1 0.18 1.17 0.19 1.63
2 0.28 1.22 0.27 1.92
4 0.48 1.39 0.50 2.04
8 1.78 2.70 0.92 2.23
16 10.07 11.12 1.68 2.96
32 / / 3.34 4.60
When performing the time test, the model is first initialized and loaded onto the GPU. Next, the start time is recorded, and the model evaluation function is executed 10 times in a loop. After each loop, the CUDA cache is manually cleared and Python’s garbage collection mechanism is triggered to ensure that memory resource is fully released. After the loop ends, the end time is recorded, and the average time for each evaluation is calculated.
@@ V. Parameter Sensitivity
In this section, we use UNet as an example to conduct a parameter sensitivity test on three hyperparameters: $M$, $U$, and $l$ in our compression algorithm, where $M$ means the maximum value of categories that can be set for all layers of the model, $U$ means the list of the number of sample nodes in the center square, and $l$ means the length of the side of center square. As shown in Table 9, it can be observed that different values of $M$, $U$, and $l$ only moderately impact the model’s performance and compression ratio, which means our algorithm is robust to hyperparameters.
@@
1 225 0.02 99.81 7.10$\times$
2 225 0.02 97.07 7.11$\times$
3 225 0.02 90.11 7.11$\times$
1 225 0.1 99.21 7.96$\times$
2 225 0.1 96.46 7.93$\times$
3 225 0.1 99.71 7.93$\times$
1 361 0.02 99.87 6.40$\times$
2 361 0.02 97.90 6.41$\times$
3 361 0.02 99.39 6.40$\times$
1 361 0.1 99.28 7.77$\times$
2 361 0.1 99.10 7.75$\times$
3 361 0.1 99.65 7.72$\times$
1 225 / 361 0.02 99.89 7.11$\times$
2 225 / 361 0.02 98.12 7.11$\times$
3 225 / 361 0.02 99.58 7.11$\times$
1 225 / 361 0.1 99.08 7.77$\times$
2 225 / 361 0.1 98.85 7.75$\times$
3 225 / 361 0.1 99.51 7.72$\times$

### v2

@@ VI. Experimental Details
  • Details of experiments related to UNet: We perform parameter compression on both UNet and the pruned UNet, and test the performance loss of the pruned models on the Carvana [1] dataset.
@@
      Pruned UNet: Specifically, the pruned UNet refers to the UNet model whose parameters are removed using the Taylor-Rank Pruning [18] with the FLOPs regularization [13] (strength=0.001). This is essentially a channel pruning method. The bar chart below illustrates the percentage of channels pruned in each layer of the pruned UNet relative to the original UNet.
@@
      Evaluation: The Dice coefficient [7], also known as the Dice similarity index, is a statistical measure used to gauge the similarity between two sets. It is particularly popular in the field of image segmentation to assess the accuracy of models by comparing the overlap between the predicted segmentation and the ground truth. Mathematically, the Dice coefficient is defined as
@@
      where $A$ represents the predicted segmentation, and $B$ is the ground truth. The value of the Dice coefficient ranges from 00 to $1$, where $1$ indicates the perfect agreement between two sets, and 00 indicates no overlap at all.
@@
      MobileNetV3: MobileNetV3 is a state-of-the-art convolutional neural network architecture designed specifically for mobile devices. Building upon the success of its predecessors, MobileNetV3 introduces novel design principles and optimization techniques to achieve a more compelling balance between accuracy and efficiency.
@@
      Dataset: The CIFAR-10 dataset [16] is a well-known benchmark in the field of machine learning and computer vision, consisting of $60,000$ $32\times 32$ colored images belonging to $10$ classes. The dataset is split into $50,000$ training images and $10,000$ testing images.
@@
  • Details of experiments related to LlaMA2-7B: We apply our compression method to the LLaMA2-7B [24] and three other performant pruned variants of the LLaMA2-7B: Sheared-LLaMA-1.3B [28], TinyLLaMA [31], and LiteLLaMA. We then examine the performance drop on these models using various metrics, including SciQ [26], WinoGrande [21], ARC-E [5], 25-shot ARC-C [5], 10-shot Hellaswag [30], 32-shot BoolQ [4], 32-shot NQ [17], 5-shot MMLU [14], and Perplexity [27] to evaluate the effects of compression by using the package lm-evaluation-harness [11].
@@
      Evaluation: In this study, most evaluation metrics for large language models (LLMs) are tested using the open-source package lm-evaluation-harness [11] which is a versatile and open-source Python package for benchmarking the ability of large language models (LLMs) across various tasks and datasets. The "($i$)" after the name of an evaluation metric means "$i$-shot".
@@ 2.1 Compression Performance
Model Compression Ratio PPL
LlaMA2-7B - 5.47
LlaMA2-7B + W8A8 SmoothQuant [29] (int8, $\alpha$=0.85) 2.00$\times$ 5.52
LlaMA2-7B + Asymmetric Quant (int8) 2.00$\times$ 5.65
LlaMA2-7B + GPTQ [9] (int4) 3.46$\times$ 5.69
LlaMA2-7B + Asymmetric Quant (int4) 4.00$\times$ 26160.34
LlaMA2-7B + HF 2.60$\times$ 5.82
Sheared-LlaMA-1.3B 2.50$\times$ 8.13
Sheared-LlaMA-1.3B + HF 12.76$\times$ 8.37
TinyLlaMA 3.06$\times$ 7.71
TinyLlaMA + HF 16.03$\times$ 7.95
LiteLlaMA 14.53$\times$ 31.82
LiteLlaMA + HF 32.05$\times$ 37.85
Figure 11: The exemplary outputs from LlaMA2-7B, Sheared-LlaMA-1.3B, TinyLlaMA, LiteLlaMA, and their compressed model by using hyper-compression. These results demonstrate that our compression method effectively preserves the models’ ability to produce meaningful and grammatically correct text.
Here, we show the compression performance of LLaMA2-7B and seven different variants based on LlaMA2-7B. First, in Table 4, we test eight important downstream tasks. By directly compressing LLaMA2-7B using our method (LLaMA2-7B + HF), a compression ratio of 2.6$\times$ is achieved with only a 0.99% decrease in the average score. When our method is combined with other LLaMA2-7B variants (Sheared-LlaMA-1.3B, TinyLlaMA, and LiteLlaMA) respectively, higher compression ratios can be achieved, with the decrease in the average score remaining within 1%, compared to the pruned ones. Moreover, Perplexity (PPL) is an important metric for evaluating the overall performance of large language models (LLMs). We conduct tests on the publicly available dataset wikitext-2-raw-v1222For more details, please visit the website of dataset wikitext-2-raw-v1: https://huggingface.co/datasets/Salesforce/wikitext.. As shown in Table 5, the PPL of LLaMA2-7B+HF increases by only 0.35 compared to the original LLaMA2-7B model, whereas the PPL increases caused by the other three LLaMA2-7B variants (Sheared-LlaMA-1.3B, TinyLlaMA, LiteLlaMA) are 2.66, 2.24, and 26.35, respectively. This demonstrates that our method not only enhances the compression ratio but also effectively maintains the model’s overall performance. We also compare the perplexity with four quantization methods: W8A8 SmoothQuant (int8, $\alpha$=0.85), Asymmetric Quant (int8), GPTQ (int4), and Asymmetric Quant (int4). Among them, both W8A8 SmoothQuant (int8, $\alpha$=0.85) and GPTQ (int4) require calibration sets for calibration during the quantization process, whereas our method does not involve calibration. Figure 11 presents two examples comparing the text outputs generated by eight different LLM models for a given prompt. The experimental data demonstrate that our method achieves a preferable compression ratio while having minimal impact on the model’s performance across various downstream tasks. Moreover, it can also integrate effectively with other compression methods such as pruning.
@@ 2.2 Compression Time
As shown in Table 6, we record how much using our method to compress eight models. The time to compress smaller models like UNet and MobileNetV3 is only 35 seconds, with Pruned-MobileNetV3 requiring only 8.83 seconds. The compression time for large language models is longer. For example, compressing LLaMA2-7B takes 53 minutes, while LiteLLaMA is compressed in just 6 minutes. The smaller the model’s size, the faster the compression is.
@@ 2.3 Inference Time
We test the inference time and the execution time of our method under different batch sizes on NVIDIA GeForce RTX 4060 and NVIDIA A40 graphics cards, respectively.
@@
1 4.35 4.43 1.50 2.06
2 3.95 4.49 1.44 2.05
4 3.63 4.77 1.45 2.22
8 3.67 7.94 1.44 2.46
When performing the time test, the model is first initialized and loaded onto the GPU. Next, the start time is recorded, and the model evaluation function is executed 10 times in a loop. After each loop, the CUDA cache is manually cleared and Python’s garbage collection mechanism is triggered to ensure that the memory resource is fully released. After the loop is finished, the end time is recorded, and the average time for each evaluation is calculated.
@@ V. Parameter Sensitivity
In this section, we use UNet as an example to conduct a parameter sensitivity test on three aforementioned hyperparameters: $M$, $U$, and $l$ in our compression algorithm, where $M$ means the maximum value of categories that can be set for all layers of the model, $U$ means the list of the number of sample nodes in the center square, and $l$ means the length of the side of the center square. As shown in Table 9, it can be observed that different values of $M$, $U$, and $l$ only moderately impact the model’s performance and compression ratio, which means our algorithm is robust to hyperparameters.
@@
1 225 0.02 99.81 7.09$\times$
2 225 0.02 97.07 7.08$\times$
3 225 0.02 90.11 7.08$\times$
1 225 0.1 99.21 7.92$\times$
2 225 0.1 96.46 7.88$\times$
3 225 0.1 99.71 7.88$\times$
1 361 0.02 99.87 6.38$\times$
2 361 0.02 97.90 6.38$\times$
3 361 0.02 99.39 6.38$\times$
1 361 0.1 99.28 7.73$\times$
2 361 0.1 99.10 7.71$\times$
3 361 0.1 99.65 7.68$\times$
1 225 / 361 0.02 99.89 7.08$\times$
2 225 / 361 0.02 98.12 7.08$\times$
3 225 / 361 0.02 99.58 7.08$\times$
1 225 / 361 0.1 99.08 7.73$\times$
2 225 / 361 0.1 98.85 7.71$\times$
3 225 / 361 0.1 99.51 7.68$\times$